# ABG-NAS: Adaptive Bayesian Genetic Neural Architecture Search for Graph Representation Learning

### v1

@@ Abstract
Effective and efficient graph representation learning is essential for enabling critical downstream tasks, such as node classification, link prediction, and subgraph search. However, existing graph neural network (GNN) architectures often struggle to adapt to diverse and complex graph structures, limiting their ability to provide robust and generalizable representations. To address this challenge, we propose ABG-NAS, a novel framework for automated graph neural network architecture search tailored for efficient graph representation learning. ABG-NAS encompasses three key components: a Comprehensive Architecture Search Space (CASS), an Adaptive Genetic Optimization Strategy (AGOS), and a Bayesian-Guided Tuning Module (BGTM). CASS systematically explores diverse propagation (P) and transformation (T) operations, enabling the discovery of GNN architectures capable of capturing intricate graph characteristics. AGOS dynamically balances exploration and exploitation, ensuring search efficiency and preserving solution diversity. BGTM further optimizes hyperparameters periodically, enhancing the scalability and robustness of the resulting architectures. Empirical evaluations on benchmark datasets (Cora, PubMed, Citeseer, and CoraFull) demonstrate that ABG-NAS consistently outperforms both manually designed GNNs and state-of-the-art neural architecture search (NAS) methods. These results highlight the potential of ABG-NAS to advance graph representation learning by providing scalable and adaptive solutions for diverse graph structures. Our code is publicly available at https://github.com/sserranw/ABG-NAS.
Index Terms:
@@
I Introduction
Graph representation learning has become essential for enabling critical tasks, such as node classification, graph classification[1], and subgraph search, by leveraging graph-structured data. Graph neural networks (GNNs) are at the forefront of this field due to their ability to aggregate and transform node features based on graph topology [2]. However, the performance of GNNs is heavily influenced by their architectural design. Traditional GNNs with fixed propagation (P) and transformation (T) components often lack the flexibility needed to adapt to diverse and complex real-world graphs, which exhibit varying node degree distributions and connectivity patterns [3]. This rigidity limits their generalizability across different tasks, and manually designing architectures for specific graphs is both time-intensive and demands substantial expertise [4]. These challenges are further compounded by the inherent diversity of graph structures, ranging from sparse graphs with low connectivity to dense graphs with highly interconnected nodes, which demand architectures capable of adapting to both extremes[5].
To address the architectural design challenges in graph representation learning, graph neural architecture search (GNAS) has emerged as a promising solution [5]. GNAS automates the process of designing GNN architectures, significantly reducing the reliance on manual intervention and enabling the discovery of architectures tailored to diverse graph structures and representation tasks. Existing GNAS methods can be broadly categorized into three types [6]: reinforcement learning (RL)-based methods [7], which optimize the search process through policy optimization; gradient-based methods, which leverage differentiable architecture search techniques; and evolutionary algorithm (EA)-based methods, including genetic algorithms (GA), particle swarm optimization (PSO), and differential evolution (DE).
RL-based methods have shown promise in GNAS by automating the search process through systematic exploration of the architecture space. Typically, GraphNAS[8] employs a recurrent neural network (RNN) to sequentially construct GNN architectures, enabling efficient exploration of the search space and providing flexible solutions for diverse datasets. Subsequently, simplified neural architecture search for graph neural networks (SNAG) [9] extended this line of research by introducing macro-level inter-layer architecture connections, offering a more comprehensive exploration of GNN designs. However, RL-based approaches are often constrained by high computational demands and fixed search spaces, which limit the diversity of generated architectures and reduce their adaptability to complex and diverse graph structures [10, 11].
@@ I Introduction
The paper is organized as follows. Section II reviews the background, related works, and preliminaries. Section III describes the ABG-NAS framework in detail, including the implementation of its main components. Section IV presents and analyzes the experimental results from multiple perspectives. Finally, Section V summarizes the findings of this paper and discusses future work.
II Related Works
II-A GNNs in Graph Representation Learning
Graph representation learning transforms graph-structured data into meaningful representations, forming the foundation for advanced methods such as graph neural networks (GNNs)[1]. GNNs, as one of the most prominent approaches, have gained attention for their ability to model intricate relationships in graphs through learnable layers[17]. Researchers have devised various GNN models with different types of layers, such as graph convolutional network (GCN) [18], graph attention network (GAT) [19], graph sample and aggregation (GraphSAGE) [20], and modality-aware graph convolutional network (MAGCN) [21], allowing for manually flexible stacking of different layers tailored to specific applications. Furthermore, Zhang et al. (2022) [14] were the first to try to decompose a GCN layer into two distinct operations: P and T. As demonstrated in Fig. 1 (a), the P operation aggregates the features from neighouring nodes to update the representation of nodes. Following this, the T operation applies a learnable weight matrix and a nonlinear activation function to enhance feature representation, allowing the GCN to capture more complex patterns.
@@ II-A GNNs in Graph Representation Learning
Previously, GCN architectures were limited to a rigid (P-T) sequence pattern, such as (P-T)-(P-T) for a 2-layer GCN and (P-T)-(P-T)-(P-T) for a 3-layer GCN, as shown in Fig. 1 (b). By decomposing layers into separate P and T operations, a wider variety of configurations, such as P-P-T-T, P-P-T, and T-T-P-T-P, can be constructed, allowing for more modular and adaptable GNN architectures.
@@ II-B EA-based GNAS
II-C Hyperparameter Optimization for GNAS
@@
$p(\mathbf{x}|y)=\begin{cases}l(\mathbf{x})&\text{if }y<y^{*}\\ g(\mathbf{x})&\text{if }y\geq y^{*}\end{cases}$ (3)
@@
III Proposed Method
In this section, we present a detailed introduction to the ABG-NAS framework. Fig. 2 depicts the overall structure of ABG-NAS, highlighting its key components and workflow. The subsequent sections will delve into the design specifics and provide corresponding pseudocode, offering a comprehensive explanation of how ABG-NAS effectively searches for and optimizes GNN architectures.
Figure 2: Overview of the ABG-NAS framework: (1) An initial population $\mathcal{Q}_{p}^{g}$ = $\mathcal{Q}_{p}^{0}$ is generated from the expanded search space; (2) Fitness evaluation is conducted for the GNN architectures in the parent population $\mathcal{Q}_{p}^{g}$, which includes the individual $I_{p}^{g*}$ with the highest fitness value; (3) The BGTM algorithm performs hyperparameter optimization in every $N_{b}$ generations thereafter; (4) The genetic operations in AGOS algorithm guides individuals go through exploration and exploitation phases based on fitness, performing selection, crossover, and mutation to generate offspring ${Q}_{o}^{g}$ from ${Q}_{p}^{g}$; (5) Offspring fitness is calculated on ${Q}_{o}^{g}$; (6) The next generation is formed by selecting individuals from both the parent population ${Q}_{p}^{g}$ and the offspring population ${Q}_{o}^{g}$ based on their fitness values. Then return to (3) to update the fitness values of the next generation ${Q}_{p}^{g}$; If $g\%N_{b}$==0, return to (3). Otherwise, if $g\leq G$, return to (4) and if $g==G$, output the best individual in ${Q}_{p}^{G}$.
III-A Problem Formulation
@@
[…]
@@
$z^{*}=\arg\max_{z\in\mathcal{Z}}\mathbb{E}_{v_{i}\in V_{val}}\psi\left[y_{i},f_{I^{*}}(\mathbf{x}_{i},\mathbf{e}_{i};z,\Theta^{*})\right]\$ (6)
@@
III-B Comprehensive Architecture Search Space
ABG-NAS introduces a more comprehensive architecture search space by expanding the diversity of P and T operations, enabling flexible combinations of components tailored for GNN neural architecture search (NAS). This design addresses the limitations of prior methods, which often restrict the search space to fixed P and T operations. It further enhances the search space by incorporating various internal configurations of the P and T operations. As illustrated in Fig. 2, eight configurations for the T operation are included, such as ReLU, Linear, and Sigmoid, thereby enabling a broader exploration of potential architectures. The expanded search space (ESS) $\mathcal{A}$ is a set containing all individual $I$ that is mathematically defined by Equation (8).
@@ III-B Comprehensive Architecture Search Space
III-C Adaptive Genetic Optimization Strategy
@@
Step 3: Adaptive Genetic Operation Switching. In genetic algorithms, selection and mating (including crossover and mutation) operations are key genetic operations that drive the evolutionary process. However, in DFG-NAS, as outlined in Section II-B Steps 3-5, these operations are applied uniformly with fixed strategies, regardless of fitness score improvements throughout the search process. This fixed approach has several drawbacks, including a higher risk of premature convergence and inefficiency in navigating the solution space due to the lack of differentiation between exploration and exploitation phases.
@@
As illustrated in Fig. 2 (4) Genetic operations in AGOS, when $\Delta\text{fitness}(g)$ is no larger than a predefined threshold $\alpha$, the algorithm enters the exploration stage; otherwise, it enters the exploitation stage. The main distinction between exploration and exploitation in AGOS lies in the values of genetic operation control factors, i.e., the selection intensity $k$, crossover probability $p_{c}$, and mutation probability $p_{m}$. During the exploration phase, higher values (denoted as $k_{1},p_{c_{1}},p_{m_{1}}$) are used to increase diversity and broaden the search space. Conversely, during the exploitation phase, lower values (denoted as $k_{2},p_{c_{2}},p_{m_{2}}$) are employed to focus on refining high-quality solutions and accelerating convergence, as demonstrated by Equation (12).
@@
Mutation. After selection and crossover, each individual in the current generation undergoes mutation with a probability $p_{m}$, which determines whether a mutation is applied. If a mutation occurs, one of four mutation types—add, remove, exchange, or alter—is randomly selected, as shown in Fig. 3 (b). Add means a random operation from $\mathcal{O}$ is inserted into a randomly selected position in the individual. For example, the individual P3-T4-P1-T2 has P2 inserted at the second position, resulting in the new individual P3-T4-P2-P1-T2. Remove represents that a randomly selected operation is deleted from the individual. For instance, in the individual P3-T4-P1-T2, the operation P1 is removed, forming the new individual P3-T4-T2. Exchange is the operation in which two randomly chosen operations within the individual are swapped. For example, in the individual P3-T4-P1-T2, the operations T4 and P1 are exchanged, resulting in the new individual P3-P1-T4-T2. Alter presents a randomly selected operation that is replaced by another operation from $\mathcal{O}$. For instance, in the individual P3-T4-P1-T2, the operation P1 is altered to T5, forming the new individual P3-T4-T5-T2.
@@
III-D Bayesian-Guided Tuning Module
@@
Both the BGTM algorithm for hyperparameter optimization and the AGOS process for G-NAS are crucial components of the proposed ABG-NAS framework, as illustrated in Fig. 2. In the rest of this section, we further present the implementation detail of BGTM and AGOS within the pseudocode of the whole ABG-NAS framework, as shown in Algorithm 1.
@@
Then, in Step 2, the optimal parent individual $I_{p}^{0*}\in\mathcal{Q}_{p}^{0}$ is identified by solving Equations (4) and (5), and the fitness scores, $\mathcal{S}_{p}^{0}$, is calculated for $\mathcal{Q}_{p}^{0}$ with the fitness function $\phi(\cdot)$, as demonstrated in Fig.2 (2) Fitness calculation.
In Step 3, a B-HPO algorithm executes to determine the optimal hyperparameter combination $z^{*}\in\mathcal{Z}$ for $I_{p}^{0*}$ by solving Equation (6), as demonstrated in Fig.2 (3) BGTM algorithm.
@@ IV Experimental Analysis
IV-A Experimental Settings
IV-A1 Datasets Details
We evaluate the proposed ABG-NAS algorithm on four widely-used public datasets: Cora[28], PubMed [29], CoraFull [30], and Citeseer[29]. These datasets differ significantly in size, sparsity, and structural complexity, enabling a comprehensive evaluation of ABG-NAS’s robustness and adaptability across diverse graph structures. Cora and PubMed are citation networks, while CoraFull is an extended version of Cora with significantly higher connectivity. In contrast, Citeseer is characterized by its higher sparsity, including 217 isolated nodes, making it particularly challenging for GNN models to extract meaningful patterns in fragmented graph data. Each dataset is split into training, validation, and testing sets in a 6:2:2 ratio. Refer to Table I for detailed dataset statistics.
TABLE I: Summary of dataset statistics used in the experiments. Din and Dout represent the average in-degree and out-degree, respectively, while Davg denotes the average degree.

### v3

@@ Abstract
Effective and efficient graph representation learning is essential for enabling critical downstream tasks, such as node classification, link prediction, and subgraph search. However, existing graph neural network (GNN) architectures often struggle to adapt to diverse and complex graph structures, limiting their ability to produce structure-aware and task-discriminative representations. To address this challenge, we propose ABG-NAS, a novel framework for automated graph neural network architecture search tailored for efficient graph representation learning. ABG-NAS encompasses three key components: a Comprehensive Architecture Search Space (CASS), an Adaptive Genetic Optimization Strategy (AGOS), and a Bayesian-Guided Tuning Module (BGTM). CASS systematically explores diverse propagation (P) and transformation (T) operations, enabling the discovery of GNN architectures capable of capturing intricate graph characteristics. AGOS dynamically balances exploration and exploitation, ensuring search efficiency and preserving solution diversity. BGTM further optimizes hyperparameters periodically, enhancing the robustness and scalability of the resulting architectures to both large-scale graphs and high-complexity models. Empirical evaluations on benchmark datasets (Cora, PubMed, Citeseer, and CoraFull) demonstrate that ABG-NAS consistently outperforms both manually designed GNNs and state-of-the-art neural architecture search (NAS) methods. These results highlight the potential of ABG-NAS to advance graph representation learning by providing adaptive solutions that scale effectively across varying graph sizes and architectural complexities. Our code is publicly available at https://github.com/sserranw/ABG-NAS.
keywords:
@@
[…]
[ltu] organization=Department of CSIT, La Trobe University, city=Melbourne, state=VIC, postcode=3086, country=Australia
[…]
[vu] organization=Institute for Sustainable Industries and Liveable Cities (ISILC), Victoria University, city=Melbourne, state=VIC, postcode=3011, country=Australia
[…]
[zjnu] organization=School of Computer Science and Technology, Zhejiang Normal University, city=Jinhua, state=Zhejiang, postcode=321004, country=China
1 Introduction
Graph representation learning has become essential for enabling critical tasks, such as node classification, graph classification, and subgraph search, by leveraging graph-structured data [1, 2]. Graph neural networks (GNNs) are at the forefront of this field due to their ability to aggregate and transform node features based on graph topology [3, 4, 5]. However, the performance of GNNs is heavily influenced by their architectural design. Traditional GNNs with fixed propagation (P) and transformation (T) components often lack the flexibility needed to adapt to diverse and complex real-world graphs, which exhibit varying node degree distributions and connectivity patterns [6]. This rigidity limits their generalizability across different tasks, and manually designing architectures for specific graphs is both time-intensive and demands substantial expertise [7]. These challenges are further compounded by the inherent diversity of graph structures, ranging from sparse graphs with low connectivity to dense graphs with highly interconnected nodes, which demand architectures capable of adapting to both extremes[8].
Graph neural architecture search (GNAS) has recently emerged as a powerful paradigm that eliminates manual trial-and-error by automatically searching for GNN architectures adaptable to a wide range of structural complexities and representation tasks [8]. It significantly reduces the reliance on manual design and enables more scalable solutions in graph representation learning. Existing GNAS methods can be broadly categorized into three types [9]: reinforcement learning (RL)-based methods [10], which optimize the search process through policy optimization; gradient-based methods, which leverage differentiable architecture search techniques; and evolutionary algorithm (EA)-based methods, including genetic algorithms (GA), particle swarm optimization (PSO), and differential evolution (DE).
RL-based methods have shown promise in GNAS by automating the search process through systematic exploration of the architecture space. Typically, GraphNAS[11] employs a recurrent neural network (RNN) to sequentially construct GNN architectures, enabling efficient exploration of the search space and providing flexible solutions for diverse datasets. Subsequently, simplified neural architecture search for graph neural networks (SNAG) [12] extended this line of research by introducing macro-level inter-layer architecture connections, offering a more comprehensive exploration of GNN designs. However, RL-based approaches are often constrained by high computational demands and fixed search spaces[13], which limit the diversity of generated architectures and reduce their adaptability to complex and diverse graph structures [14, 15, 16].
@@ 1 Introduction
The paper is organized as follows. Section 2 reviews the background, related works, and preliminaries. Section 3 describes the ABG-NAS framework in detail, including the implementation of its main components. Section 4 presents and analyzes the experimental results from multiple perspectives. Finally, Section 5 summarizes the findings of this paper and discusses future work.
2 Related Works
2.1 GNNs in Graph Representation Learning
Graph representation learning transforms graph-structured data into meaningful representations, forming the foundation for advanced methods such as graph neural networks (GNNs)[1, 4]. GNNs, as one of the most prominent approaches, have gained attention for their ability to model intricate relationships in graphs through learnable layers[22, 23]. Researchers have devised various GNN models with different types of layers, such as graph convolutional network (GCN) [24], graph attention network (GAT) [25], graph sample and aggregation (GraphSAGE) [26], and modality-aware graph convolutional network (MAGCN) [27], allowing for manually flexible stacking of different layers tailored to specific applications. Furthermore, Zhang et al. (2022) [19] were the first to try to decompose a GCN layer into two distinct operations: P and T. As demonstrated in Figure 1 (a), the P operation aggregates the features from neighouring nodes to update the representation of nodes. Following this, the T operation applies a learnable weight matrix and a nonlinear activation function to enhance feature representation, allowing the GCN to capture more complex patterns.
@@ 2.1 GNNs in Graph Representation Learning
Previously, GCN architectures were limited to a rigid (P-T) sequence pattern, such as (P-T)-(P-T) for a 2-layer GCN and (P-T)-(P-T)-(P-T) for a 3-layer GCN, as shown in Figure 1 (b). By decomposing layers into separate P and T operations, a wider variety of configurations, such as P-P-T-T, P-P-T, and T-T-P-T-P, can be constructed, allowing for more modular and adaptable GNN architectures.
@@ 2.2 EA-based GNAS
2.3 Hyperparameter Optimization for GNAS
@@
[…]
@@
3 Proposed Method
In this section, we present a detailed introduction to the ABG-NAS framework. Figure 2 depicts the overall structure of ABG-NAS, highlighting its key components and workflow. The subsequent sections will delve into the design specifics and provide corresponding pseudocode, offering a comprehensive explanation of how ABG-NAS effectively searches for and optimizes GNN architectures.
Figure 2: Overview of the ABG-NAS framework: (1) An initial population $\mathcal{Q}_{p}^{g}$ = $\mathcal{Q}_{p}^{0}$ is generated from the expanded search space; (2) Fitness evaluation is conducted for the GNN architectures in the parent population $\mathcal{Q}_{p}^{g}$, which includes the individual $I_{p}^{g*}$ with the highest fitness value; (3) The BGTM algorithm performs hyperparameter optimization in every $N_{b}$ generations thereafter; (4) The genetic operations in AGOS algorithm guide individuals go through exploration and exploitation phases based on fitness, performing selection, crossover, and mutation to generate offspring ${Q}_{o}^{g}$ from ${Q}_{p}^{g}$; (5) Offspring fitness is calculated on ${Q}_{o}^{g}$; (6) The next generation is formed by selecting individuals from both the parent population ${Q}_{p}^{g}$ and the offspring population ${Q}_{o}^{g}$ based on their fitness values. Then return to (3) to update the fitness values of the next generation ${Q}_{p}^{g}$; If $g\%N_{b}$==0, return to (3). Otherwise, if $g\leq G$, return to (4) and if $g==G$, output the best individual in ${Q}_{p}^{G}$.
3.1 Problem Formulation
@@
$\text{s.t.}\quad\Theta^{*}$ $=\arg\min_{\Theta}\mathbb{E}_{v_{i}\in V_{train}}\ell\left[y_{i},f_{I}(\mathbf{x}_{i},\mathbf{e}_{i};z,\Theta)\right].$ (5)
@@
$z^{*}=\arg\max_{z\in\mathcal{Z}}\mathbb{E}_{v_{i}\in V_{val}}\psi\left[y_{i},f_{I^{*}}(\mathbf{x}_{i},\mathbf{e}_{i};z,\Theta^{*})\right]\\$ (6)
@@
3.2 Comprehensive Architecture Search Space
ABG-NAS introduces a more comprehensive architecture search space by expanding the diversity of P and T operations, enabling flexible combinations of components tailored for GNN neural architecture search (NAS). This design addresses the limitations of prior methods, which often restrict the search space to fixed P and T operations. It further enhances the search space by incorporating various internal configurations of the P and T operations. As illustrated in Figure 2, eight configurations for the T operation are included, such as ReLU, Linear, and Sigmoid, thereby enabling a broader exploration of potential architectures. The expanded search space (ESS) $\mathcal{A}$ is a set containing all individual $I$ that is mathematically defined by Equation (8).
@@ 3.2 Comprehensive Architecture Search Space
3.3 Adaptive Genetic Optimization Strategy
@@
Step 3: Adaptive Genetic Operation Switching. In genetic algorithms, selection and mating (including crossover and mutation) operations are key genetic operations that drive the evolutionary process. However, in DFG-NAS, as outlined in Section 2.2 Steps 3-5, these operations are applied uniformly with fixed strategies, regardless of fitness score improvements throughout the search process. This fixed approach has several drawbacks, including a higher risk of premature convergence and inefficiency in navigating the solution space due to the lack of differentiation between exploration and exploitation phases.
@@
As illustrated in Figure 2 (4) Genetic operations in AGOS, when $\Delta\text{fitness}(g)$ is no larger than a predefined threshold $\alpha$, the algorithm enters the exploration stage; otherwise, it enters the exploitation stage. The main distinction between exploration and exploitation in AGOS lies in the values of genetic operation control factors, i.e., the selection intensity $k$, crossover probability $p_{c}$, and mutation probability $p_{m}$. During the exploration phase, higher values (denoted as $k_{1},p_{c_{1}},p_{m_{1}}$) are used to increase diversity and broaden the search space. Conversely, during the exploitation phase, lower values (denoted as $k_{2},p_{c_{2}},p_{m_{2}}$) are employed to focus on refining high-quality solutions and accelerating convergence, as demonstrated by Equation (12).
@@
Mutation. After selection and crossover, each individual in the current generation undergoes mutation with a probability $p_{m}$, which determines whether a mutation is applied. If a mutation occurs, one of four mutation types—add, remove, exchange, or alter—is randomly selected, as shown in Figure 3 (b). Add means a random operation from $\mathcal{O}$ is inserted into a randomly selected position in the individual. For example, the individual P3-T4-P1-T2 has P2 inserted at the second position, resulting in the new individual P3-T4-P2-P1-T2. Remove represents that a randomly selected operation is deleted from the individual. For instance, in the individual P3-T4-P1-T2, the operation P1 is removed, forming the new individual P3-T4-T2. Exchange is the operation in which two randomly chosen operations within the individual are swapped. For example, in the individual P3-T4-P1-T2, the operations T4 and P1 are exchanged, resulting in the new individual P3-P1-T4-T2. Alter presents a randomly selected operation that is replaced by another operation from $\mathcal{O}$. For instance, in the individual P3-T4-P1-T2, the operation P1 is altered to T5, forming the new individual P3-T4-T5-T2.
@@
3.4 Bayesian-Guided Tuning Module
@@
Both the BGTM algorithm for hyperparameter optimization and the AGOS process for G-NAS are crucial components of the proposed ABG-NAS framework, as illustrated in Figure 2. In the rest of this section, we further present the implementation detail of BGTM and AGOS within the pseudocode of the whole ABG-NAS framework, as shown in Algorithm 1.
@@
Then, in Step 2, the optimal parent individual $I_{p}^{0*}\in\mathcal{Q}_{p}^{0}$ is identified by solving Equations (4) and (5), and the fitness scores, $\mathcal{S}_{p}^{0}$, is calculated for $\mathcal{Q}_{p}^{0}$ with the fitness function $\phi(\cdot)$, as demonstrated in Figure2 (2) Fitness calculation.
In Step 3, a B-HPO algorithm executes to determine the optimal hyperparameter combination $z^{*}\in\mathcal{Z}$ for $I_{p}^{0*}$ by solving Equation (6), as demonstrated in Figure2 (3) BGTM algorithm.
@@ 4 Experimental Analysis
4.1 Experimental Settings
4.1.1 Datasets Details
We evaluate the proposed ABG-NAS algorithm on four widely-used public datasets: Cora[34], PubMed [35], CoraFull [36], and Citeseer[35]. These datasets differ significantly in size, sparsity, and structural complexity, enabling a comprehensive evaluation of ABG-NAS’s robustness and adaptability across diverse graph structures. Cora and PubMed are citation networks, while CoraFull is an extended version of Cora with significantly higher connectivity. In contrast, Citeseer is characterized by its higher sparsity, including 217 isolated nodes, making it particularly challenging for GNN models to extract meaningful patterns in fragmented graph data. Each dataset is split into training, validation, and testing sets in a 6:2:2 ratio. Refer to Table 1 for detailed dataset statistics.
Table 1: Summary of dataset statistics used in the experiments. Din and Dout represent the average in-degree and out-degree, respectively, while Davg denotes the average degree.